# The Hidden Attention of Mamba Models

### v1

@@ Abstract
The Mamba layer offers an efficient selective state space model (SSM) that is highly effective in modeling multiple domains including NLP, long-range sequences processing, and computer vision. Selective SSMs are viewed as dual models, in which one trains in parallel on the entire sequence via IO-aware parallel scan, and deploys in an autoregressive manner. We add a third view and show that such models can be viewed as attention-driven models. This new perspective enables us to compare the underlying mechanisms to that of the self-attention layers in transformers and allows us to peer inside the inner workings of the Mamba model with explainability methods. Our code is publicly available [^1] https://github.com/AmeenAli/HiddenMambaAttn.
@@ 1 Introduction
Despite their growing success, the information-flow dynamics between tokens in Mamba models and the way they learn remain largely unexplored. Critical questions about their learning mechanisms, particularly how they capture dependencies and their resemblance to other established layers like RNNs, CNNs, or attention mechanisms, remain unanswered. Additionally, the lack of interoperability methods for these models may pose a significant hurdle to debugging them and may also reduce their applicability in socially sensitive domains in which explainability is required.
Motivated by these gaps, our research aims to provide insights into the dynamics of the Mamba models and develop methodologies for their interpretation. While the traditional views of state-space models are through the lens of convolutional or recurrent layers [25], we show that selective state-space layers are a form of attention models. This is achieved through a novel reformulation of Mamba computation using a data-control linear operator, unveiling hidden attention matrices within the Mamba layer. This enables us to employ well-established interpretability and explainability techniques, commonly used in transformer realms, to devise the first set of tools for interpreting Mamba models. Furthermore, our analysis of implicit attention matrices offers a direct framework for comparing the properties and inner-representations of transformers [53] and selective-state space layers.
Our main contributions encompass the following main aspects: (i) We shed light on the fundamental nature of Mamba models, by showing that they rely on implicit attention, which is implemented by a unique data-control linear operator, as illustrated in Fig. 1 (right). (ii) Our analysis reveals that Mamba models give rise to three orders of magnitude more attention matrices than transformers. (iii) We provide a set of explainability and interpretability tools based on these hidden attention matrices. (iv) In the domain of computer vision, for comparable model sizes, Mamba model-based attention shows comparable explainability metrics results to that of transformers.
Figure 1: Three Perspectives of the Selective State-Space Layer:(Left) Selective State-Space Models (SSMs) can be efficiently computed with linear complexity using parallel scans, allowing for effective parallelization on modern hardware, such as GPUs. (Middle) Similar to SSMs, the selective state-space layer can be computed via a time-variant recurrent rule. (Right) A new view of the selective SSM layer, showing that it uses attention similarly to transformers (see Eq. 12). Our view enables the generation of attention maps, offering valuable applications in areas like XAI.
@@ 2 Background
where $Q$, $K$, and $V$ represent queries, keys, and values, respectively, and $d_{k}$ is the dimension of the keys. Additionally, the Transformer utilizes $H$ attention heads to parallelly process information, allowing the model to capture various dependencies. The attention matrix $\alpha$ enables the models to weigh the importance of tokens based on their contribution to the context, and they can also used for interpretability [7], explainability [13], and improved classification [52, 14].
@@
Since the recurrence is linear, Eq. 3 can also be expressed as a convolution, via a convolution kernel $K:=(k_{1},\cdots,k_{L})$, where $k_{i}=C\bar{A}^{i-1}\bar{B}$. Thus, allowing sub-quadratic complexity in sequence length. The equivalence between the recurrent and convolution provides a versatile framework that enables parallel and efficient training with sub-quadratic complexity with the convolution view, alongside a faster recurrent view, facilitating the acceleration of autoregressive generation by decoupling step complexity from sequence length. As the layer defined as a map from $\mathbb{R^{L}}$ to $\mathbb{R^{L}}$, to processes $D$ channels the layer employ $D$ independent copies of itself.
S6 Layers A recent development in state space layers is selective SSMs [23] (S6), which show outstanding performance in NLP [4, 44, 56], vision [33, 64], graph classification [54, 9], and more. These models rely on time-variant state space layers, namely, the discrete matrices $\bar{A},\bar{B},$ and $C$ of each channel are modified over the $L$ time steps depending on the input sequence. As opposed to traditional state-space layers, which operate individually on each channel, selective state-space layers compute the SSM matrices $\bar{A}_{i},\bar{B}_{i},C_{i}$ for all $i\leq L$ based on all the channels, and then apply the time-variant recurrent rule individually for each channel. Hence, we denote the enitre input sequence by $\hat{x}:=(\hat{x}_{1},\cdots,\hat{x}_{L})\in\mathbb{R}^{L\times D}$ where $\hat{x}_{i}\in\mathbb{R}^{D}$. The per-time discrete matrices $\bar{A_{i}},\bar{B_{i}},$ and $C_{i}$ are defined as follows:
$B_{i}=S_{B}\hat{x}_{i},\quad C_{i}=S_{C}\hat{x}_{i},\quad\Delta_{i}=\text{softplus}(S_{\Delta}\hat{x}_{i})$ (4)
$f_{A}(\Delta_{i},A)=\exp(\Delta_{i}A),\quad f_{B}(\Delta_{i},A,B_{i})=\Delta_{i}B$ (5)
@@
where $f_{A},f_{B}$ represents the discretization rule, $S_{A},S_{B},S_{\Delta}$ are linear projection layers, and SoftPlus is an element-wise function that is a smooth approximation of ReLU. While previous state-space layers employ complex-valued SSMs and non-diagonal matrices, Mamba employs real-diagonal parametrization for the system matrices.
The motivation for input-dependent time-variant layers is to make those recurrent layers more expressive and flexible, allowing them to capture more complex dependencies. While other input-dependent time-variant mechanisms have been proposed in previous works through gated RNNs, S5 layer [49], or adaptive filtering via input-dependent IIR filters [36], Mamba significantly improves on these layers by presenting a flexible, yet still efficient, approach. This efficiency was achieved via the IO-aware implementation of associative scans, which can be parallelized on modern hardware via work-efficient parallel scanners [11, 39].
Mamba The Mamba block is built on top of the selective state-space layer, Conv1D and other elementwize operators. Inspired by the architecure of Gated MLP and H3 [19], and given an input $\hat{x}^{\prime}:=(\hat{x}^{\prime}_{1},\cdots\hat{x}^{\prime}_{L})$ it is defined as follows:
@@
$\hat{y}^{\prime}=\text{Linear}(\text{Selective SSM}(\hat{x^{\prime}})\otimes\hat{z})),\quad\hat{y}=\hat{y}^{\prime}+\hat{x}^{\prime},\quad\ddot{y}=\text{LayerNorm}(\hat{y})$ (8)
where $\otimes$ is elementwise multiplication. Mamba models contains $\Lambda$ stacked mamba blocks and $D$ channels per block, and we denote the tensors in the i-th block and j-th channel with a superscript, where the first index refers to the block number.
@@
The contributions most closely aligned with ours are those specifically tailored for transformer explainability. Abnar and Zuidema [1] introduce the Attention-Rollout method, which aggregates the attention matrices across different layers by analyzing the paths in the inter-layer pairwise attention graph. Chefer et al. [13, 12] combine LRP scores [6] with the attention gradients in order to obtain a class-specific relevance scores. Ali et al. [3] enhanced attributions by treating the non-linear Softmax and LayerNorm operators as a constant, thereby attributing relevance exclusively through the value path, disregarding these operators Yuan et al. [63] treats the output token representations as states in a Markov chain in which the transition matrix is built using attention weights.
@@ 3.1 Hidden Attention Matrices In Selective State Spaces Layers
By considering the initial conditions $h_{0}=0$, unroling Eq. 2 for one channel yields:
$h_{1}=\bar{B}_{1}x_{1},\quad y_{1}=C_{1}\bar{B}_{1}x_{1}$ (9)
$h_{2}=\bar{A}_{2}\bar{B}_{1}x_{1}+\bar{B}_{2}x_{2},\quad y_{2}=C_{2}\bar{A}_{2}\bar{B}_{1}x_{1}+C_{2}\bar{B}_{2}x_{2}$ (10)
@@
Eq. 12 and 13 links $\tilde{\alpha}$ to the conventional standard attention matrix (Eq. 1), and highlights that S6 can be considered a variant of causal self-attention.
@@
where $C_{i}[m],A_{k}[m,m],B_{j}[m]\in\mathbb{R}$, plugging it into Eq. 13 yiels:
@@
Note that while equations 11, and 13 contains matrix multiplication, Eq. 14 relies on element-wise multiplication.
An interesting observation arising from Eq. 15 is that a single channel of S6 produces $N$ inner attention matrices, which can be computed by:
$\forall m\in[N]:C_{i}[m]\Big{(}\Pi_{k=j+1}^{i}\bar{A}_{k}[m,m]\Big{)}\bar{B}_{j}[m]$ (16)
These $N$ per-coordinate attention matrices are aggregated in Eq. 15 into a single global attention matrix $\tilde{\alpha}$, per channel. However, in the Transformer, a single attention matrix is produced by each of the $H$ attention heads. Given that the number of channels in Mamba models $D$ is typically a hundred times greater than the number of heads in a transformer (for example, Vision-Mamba-Tiny has $D=384$ channels, compared to $H=3$ heads in DeiT-Tiny), the Mamba layer generates approximately $\frac{DN}{H}\approx 100N$ more attention matrices than the original self-attention layer.
To further understand the structure and characterization of these attention matrices, we will express those hidden attention matrices $\tilde{\alpha}$ for each channel $d$ as a direct function of the input $\hat{x}$. To to so, we first substitute Eq.4,5 and Eq.6 into Eq. 13, and obtain:
$\tilde{\alpha}_{i,j}=S_{C}\hat{x}_{i}\Big{(}\Pi_{k=j+1}^{i}\exp\Big{(}\text{softplus}(S_{\Delta}\hat{x}_{k})A\Big{)}\Big{)}\text{softplus}(S_{\Delta}\hat{x}_{j})S_{B}\hat{x}_{j}=$ (17)
$S_{C}\hat{x}_{i}\Big{(}\exp\ \Big{(}\sum_{k=j+1}^{i}\text{softplus}(S_{\Delta}\hat{x}_{k})\Big{)}A\Big{)}\text{softplus}(S_{\Delta}\hat{x}_{j})S_{B}\hat{x}_{j}$ (18)
@@
$\tilde{\alpha}_{i,j}\approx S_{C}\hat{x}_{i}\left(\exp\ \left(\sum_{\begin{subarray}{c}k=j+1\\ S_{\Delta}\hat{x}_{k}>0\end{subarray}}^{i}(S_{\Delta}\hat{x}_{k})\right)A\right)\text{ReLU}(S_{\Delta}\hat{x}_{j})S_{B}\hat{x}_{j}$ (19)
@@
$\tilde{Q_{i}}:=S_{C}\hat{x}_{i},\textbf{ }\tilde{K_{j}}:=\text{ReLU}(S_{\Delta}\hat{x}_{j})S_{B}\hat{x}_{j},\textbf{ }\tilde{H}_{i,j}:=\exp\Big{(}\sum_{\begin{subarray}{c}k=j+1\\ S_{\Delta}\hat{x}_{k}>0\end{subarray}}^{i}(S_{\Delta}\hat{x}_{k})\Big{)}A$ (20)
@@
$\tilde{\alpha}_{i,j}\approx\tilde{Q_{i}}\tilde{K_{j}}\tilde{H}_{i,j}$ (21)
@@
This distinction between the self-attention and Mamba, captured by $\tilde{H}_{i,j}$ could be a key factor in enabling Mamba-based models to understand and utilize continuous historical context within sequences more efficiently than Transforme models.
[…]
A distinct divergence between Mamba’s attention mechanism and traditional self-attention lies in the latter’s utilization of a per-row softmax function. It’s essential to recognize that various attention models have either omitted the softmax [35] or substituted it with element-wise neural activations [28, 59, 38, 65], achieving comparable outcomes to the original framework.
@@ 3.2 Application to Attention Rollout
As our class-agnostic explainability teqchniqe for Mamba models, we built our method on top of the Attention-Rollout [1] method. For simplicity, we assume that we deal with a vision mamba model, which operates on sequences of size $L+1$, where $L$ is the sequence length obtained from the $\sqrt{L}\times\sqrt{L}$ image patcehs, with a classification (CLS) token appended to the sequence’s end.
@@
$\forall\lambda\in[\Lambda]:\quad\tilde{\mathbb{\alpha}}^{\lambda}=\mathbb{I}_{L+1}+\mathop{\mathbb{E}}_{d\in[D]}(\tilde{\alpha}^{\lambda,d})$ (22)
where ${\tilde{\alpha}}^{\lambda}\in\mathbb{R}^{(L+1)\times(L+1)}$ and $\mathbb{I}_{L+1}\in\mathbb{R}^{(L+1)\times(L+1)}$ is an identity matrix utilized to incorporate the influence of skip connections along the layers.
@@
Note that each row of $\rho$ corresponds to a relevance map for each token given the other tokens. In the context of this study, which concentrates on classification models, our attention analysis exclusively directs attention to the CLS token, thus, we derive the final relevance map from the raw associated with the CLS token in the output matrix, denoted by $\rho_{\text{CLS}}\in\mathbb{R}^{L}$, which contains the relevance scores evaluating each token’s influence on the classification token. Finally, to get the final explanation heatmap we reshape $\rho_{\text{CLS}}\in\mathbb{R}^{L}$ to $\sqrt{L}\times\sqrt{L}$ and upsample it back to the size of the original image using bilinear interpolation.
Although Mamba models are causal by defintion, resulting in causal hidden attention matrices, our method can be straightforwardly extended to a bidirectional setting. This adaptation involves modifying Eq. 22 so that $\tilde{\mathbb{\alpha}}^{\lambda,d}$ becomes the outcome of summing the (two) per-direction matrices of the $\lambda$-layer and the $d$-channel.

### v2

@@ Abstract
The Mamba layer offers an efficient selective state space model (SSM) that is highly effective in modeling multiple domains, including NLP, long-range sequence processing, and computer vision. Selective SSMs are viewed as dual models, in which one trains in parallel on the entire sequence via an IO-aware parallel scan, and deploys in an autoregressive manner. We add a third view and show that such models can be viewed as attention-driven models. This new perspective enables us to empirically and theoretically compare the underlying mechanisms to that of the self-attention layers in transformers and allows us to peer inside the inner workings of the Mamba model with explainability methods. Our code is publicly available [^1] https://github.com/AmeenAli/HiddenMambaAttn.
@@ 1 Introduction
Despite their growing success, the information-flow dynamics between tokens in Mamba models and the way they learn remain largely unexplored. Critical questions about their learning mechanisms, particularly how they capture dependencies and their resemblance to other established layers, such as RNNs, CNNs, or attention mechanisms, remain unanswered. Additionally, the lack of interoperability methods for these models may pose a significant hurdle to debugging them and may also reduce their applicability in socially sensitive domains in which explainability is required.
Motivated by these gaps, our research aims to provide insights into the dynamics of Mamba models and develop methodologies for their interpretation. While the traditional views of state-space models are through the lens of convolutional or recurrent layers [32], we show that selective state-space layers are a form of attention models. This is achieved through a novel reformulation of Mamba computation using a data-control linear operator, unveiling hidden attention matrices within the Mamba layer. This enables us to employ well-established interpretability and explainability techniques, commonly used in transformer realms, to devise the first set of tools for interpreting Mamba models. Furthermore, our analysis of implicit attention matrices offers a direct framework for comparing the properties and inner representations of transformers [70] and Mamba models.
Our main contributions encompass the following main aspects: (i) We shed light on the fundamental nature of Mamba models, by showing that they rely on implicit attention, which is implemented by a unique data-control linear operator, as illustrated in Fig. 1 (right). (ii) Our analysis reveals that Mamba models give rise to three orders of magnitude more attention matrices than transformers. (iii) We provide a set of explainability and interpretability tools based on these hidden attention matrices. (iv) For comparable model sizes, Mamba model-based attention shows comparable explainability metrics results to that of transformers. (v) We present a theoretical analysis of the evolution of attention capabilities in state-space models and their expressiveness, offering a deeper understanding of the factors that contribute to Mamba’s effectiveness.
Figure 1: Three Perspectives of the Selective State-Space Layer:(Left) Selective State-Space Models (SSMs) can be efficiently computed with linear complexity using parallel scans, allowing for effective parallelization on modern hardware, such as GPUs. (Middle) Similar to SSMs, the selective state-space layer can be computed via a time-variant recurrent rule. (Right) A new view of the selective SSM layer, showing that it uses attention similarly to transformers (see Eq. 11). Our view enables the generation of attention maps, offering valuable applications in areas such as XAI.
@@ 2 Background
where $Q$, $K$, and $V$ represent queries, keys, and values, respectively, and $d_{k}$ is the dimension of the keys. Additionally, the Transformer utilizes $H$ attention heads to process information in parallel, allowing the model to capture various dependencies. The attention matrix $\alpha$ enables the models to weigh the importance of tokens based on their contribution to the context, and they can also used for interpretability [8], explainability [15], and improved classification [69, 16].
@@
Since the recurrence is linear, Eq. 3 can also be expressed as a convolution, via a convolution kernel $K:=(k_{1},\cdots,k_{L})$, where $k_{i}=C\bar{A}^{i-1}\bar{B}$, thus allowing sub-quadratic complexity in sequence length. The equivalence between the recurrence and the convolution provides a versatile framework that enables parallel and efficient training with sub-quadratic complexity with the convolution view, alongside a faster recurrent view, facilitating the acceleration of autoregressive generation by decoupling step complexity from sequence length. As the layer defined as a map from $\mathbb{R^{L}}$ to $\mathbb{R^{L}}$, to process $D$ channels the layer employs $D$ independent copies of itself.
S6 Layers A recent development in state space layers is selective SSMs [30] (S6), which show outstanding performance in NLP [5, 58, 73], vision [44, 82], graph classification [71, 10], and more. These models rely on time-variant SSMs, namely, the discrete matrices $\bar{A},\bar{B},$ and $C$ of each channel are modified over the $L$ time steps depending on the input sequence. As opposed to traditional state-space layers, which operate individually on each channel, selective state-space layers compute the SSM matrices $\bar{A}_{i},\bar{B}_{i},C_{i}$ for all $i\leq L$ based on all the channels, and then apply the time-variant recurrent rule individually for each channel. Hence, we denote the entire input sequence by $\hat{x}:=(\hat{x}_{1},\cdots,\hat{x}_{L})\in\mathbb{R}^{L\times D}$ where $\hat{x}_{i}\in\mathbb{R}^{D}$. The per-time discrete matrices $\bar{A_{i}},\bar{B_{i}},$ and $C_{i}$ are defined as follows:
$B_{i}=S_{B}(\hat{x}_{i}),\quad C_{i}=S_{C}(\hat{x}_{i}),\quad\Delta_{i}=\text{softplus}(S_{\Delta}(\hat{x}_{i}))$ (4)
$f_{A}(\Delta_{i},A)=\exp(\Delta_{i}A),\quad f_{B}(\Delta_{i},A,B_{i})=\Delta_{i}B_{i},$ (5)
@@
where $f_{A},f_{B}$ represents the discretization rule, $S_{B},S_{C},S_{\Delta}$ are linear projection layers, and SoftPlus is an elementwise function that is a smooth approximation of ReLU. While previous state-space layers employ complex-valued SSMs and non-diagonal matrices, Mamba employs real-diagonal parametrization.
The motivation for input-dependent time-variant layers is to make those recurrent layers more expressive and flexible, allowing them to capture more complex dependencies. While other input-dependent time-variant mechanisms have been proposed in previous works through gated RNNs, the S5 layer [66], or adaptive filtering via input-dependent IIR filters [47], Mamba significantly improves on these layers by presenting a flexible, yet still efficient, approach. This efficiency was achieved via the IO-aware implementation of associative scans, which can be parallelized on modern hardware via work-efficient parallel scanners [12, 50].
Mamba The Mamba block is built on top of the selective state-space layer, Conv1D and other elementwise operators. Inspired by the architecure of Gated MLP and H3 [24], and given an input $\hat{x}^{\prime}:=(\hat{x}^{\prime}_{1},\cdots\hat{x}^{\prime}_{L})$ it is defined as follows:
@@
$\hat{y}^{\prime}=\text{Linear}(\text{Selective SSM}(\hat{x})\otimes\hat{z})),\quad\hat{y}=\text{LayerNorm}(\hat{y}^{\prime}+\hat{x}^{\prime})$ (8)
where $\otimes$ is elementwise multiplication. Mamba models contain $\Lambda$ stacked mamba blocks and $D$ channels per block, and we denote the tensors in the i-th block and j-th channel with a superscript, where the first index refers to the block number.
@@
The contributions most closely aligned with ours are those specifically tailored for transformer explainability. Abnar and Zuidema [1] introduce the Attention-Rollout method, which aggregates the attention matrices across different layers by analyzing the paths in the inter-layer pairwise attention graph. Chefer et al. [15, 14] combine LRP scores [7] with the attention gradients to obtain class-specific relevance scores. Ali et al. [4] enhanced attributions by treating the non-linear Softmax and LayerNorm operators as a constant, thereby attributing relevance exclusively through the value path, disregarding these operators. Yuan et al. [81] treats the output token representations as states in a Markov chain in which the transition matrix is built using attention weights.
@@ 3.1 Hidden Attention Matrices In Selective State Spaces Layers
By considering the initial conditions $h_{0}=0$, unrolling Eq. 2 yields:
$h_{1}=\bar{B}_{1}x_{1},\quad y_{1}=C_{1}\bar{B}_{1}x_{1},\quad h_{2}=\bar{A}_{2}\bar{B}_{1}x_{1}+\bar{B}_{2}x_{2},\quad y_{2}=C_{2}\bar{A}_{2}\bar{B}_{1}x_{1}+C_{2}\bar{B}_{2}x_{2}$ (9)
@@
Eq. 11 and 12 link $\tilde{\alpha}$ to the conventional standard attention matrix (Eq. 1), and highlight that S6 can be considered a variant of causal self-attention.
@@
where $C_{i}[m],A_{k}[m,m],B_{j}[m]\in\mathbb{R}$, plugging it into Eq. 12 yields:
@@
Note that while equations 10 and 12 contain matrix multiplication, Eq. 13 relies on elementwise multiplication.
An interesting observation arising from Eq. 14 is that a single channel of S6 produces $N$ inner attention matrices $C_{i}[m]\Big{(}\Pi_{k=j+1}^{i}\bar{A}_{k}[m,m]\Big{)}\bar{B}_{j}[m]$, which are summed up over $m$ to obtain $\tilde{\alpha}$. In contrast, in the Transformer, a single attention matrix is produced by each of the $H$ attention heads. Given that the number of channels in Mamba models $D$ is typically a hundred times greater than the number of heads in a transformer (for example, Vision-Mamba-Tiny has $D=384$ channels, compared to $H=3$ heads in DeiT-Tiny), the Mamba layer generates approximately $\frac{DN}{H}\approx 100N$ more attention matrices than the original self-attention layer.
To further understand the structure and characterization of these attention matrices, we will express the hidden attention matrices $\tilde{\alpha}$ for each channel $d$ as a direct function of the input $\hat{x}$. To do so, we first substitute Eq.45 and Eq.6 into Eq. 12, and obtain:
$\tilde{\alpha}_{i,j}=S_{C}(\hat{x}_{i})\Big{(}\Pi_{k=j+1}^{i}\exp\Big{(}\text{softplus}(S_{\Delta}(\hat{x}_{k}))A\Big{)}\Big{)}\text{softplus}(S_{\Delta}(\hat{x}_{j}))S_{B}(\hat{x}_{j})=$ (15)
$S_{C}(\hat{x}_{i})\Big{(}\exp\ \Big{(}\sum_{k=j+1}^{i}\text{softplus}(S_{\Delta}(\hat{x}_{k}))\Big{)}A\Big{)}\text{softplus}(S_{\Delta}(\hat{x}_{j}))S_{B}(\hat{x}_{j})$ (16)
@@
$\tilde{\alpha}_{i,j}\approx S_{C}(\hat{x}_{i})\left(\exp\ \left(\sum_{\begin{subarray}{c}k=j+1\\ S_{\Delta}(\hat{x}_{k})>0\end{subarray}}^{i}S_{\Delta}(\hat{x}_{k})\right)A\right)\text{ReLU}(S_{\Delta}(\hat{x}_{j}))S_{B}(\hat{x}_{j})$ (17)
@@
$\tilde{Q_{i}}:=S_{C}(\hat{x}_{i}),\textbf{ }\tilde{K_{j}}:=\text{ReLU}(S_{\Delta}(\hat{x}_{j})S_{B}(\hat{x}_{j}),\textbf{ }\tilde{H}_{i,j}:=\exp\Big{(}\sum_{\begin{subarray}{c}k=j+1\\ S_{\Delta}(\hat{x}_{k})>0\end{subarray}}^{i}S_{\Delta}(\hat{x}_{k})\Big{)}A$ (18)
@@
$\tilde{\alpha}_{i,j}\approx\tilde{Q_{i}}\tilde{H}_{i,j}\tilde{K_{j}}$ (19)
@@
This distinction between self-attention and Mamba, captured by $\tilde{H}_{i,j}$ could be a key factor in enabling Mamba models to understand and utilize continuous historical context within sequences more efficiently than attention.
Furthermore, Eq. 19, and 18 offer further insights into the characterization of the hidden attention matrices by demonstrating that the only terms modified across channels are $A$ and $\Delta_{i}$, which influence the values of $\tilde{H}_{i,j}$ and $\tilde{K}_{j}$ through the discretization rule in Eq. 5. Hence, all the hidden attention matrices follow a common pattern, distinguished by the keys $\tilde{K}_{j}$ via $\Delta_{i}$ and the significance of the history $\tilde{H}_{i,j}$ via $A$ and $\Delta_{i}$.
A distinct divergence between Mamba’s attention mechanism and traditional self-attention lies in the latter’s utilization of a per-row softmax function. It is essential to recognize that various attention models have either omitted the softmax [46] or substituted it with elementwise neural activations [38, 77, 49, 83], achieving comparable outcomes to the original framework.
The softmax operator is known to lead to oversmoothing [3, 74]. As we show in the Appendix 0.A, Mamba attention layers have lower inter-token smoothing than Transformer attention.
@@ 3.2 Application to Attention Rollout
As our class-agnostic explainability teqchniqe for Mamba models, we built our method on top of the Attention-Rollout [1] method. For simplicity, we assume that we are dealing with a vision mamba model, which operates on sequences of size $L+1$, where $L$ is the sequence length obtained from the $\sqrt{L}\times\sqrt{L}$ image patches, with a classification (CLS) token appended to the end of the sequence.
@@
$\forall\lambda\in[\Lambda]:\quad\tilde{\mathbb{\alpha}}^{\lambda}=\mathbb{I}_{L+1}+\mathop{\mathbb{E}}_{d\in[D]}(\tilde{\alpha}^{\lambda,d}),\quad{\tilde{\alpha}}^{\lambda}\in\mathbb{R}^{(L+1)\times(L+1)}$ (20)
where $\mathbb{I}_{L+1}\in\mathbb{R}^{(L+1)\times(L+1)}$ is an identity matrix utilized to incorporate the influence of skip connections along the layers.
@@
Note that each row of $\rho$ corresponds to a relevance map for each token, given the other tokens. In the context of this study, which concentrates on classification models, our attention analysis directs attention exclusively to the CLS token. Thus, we derive the final relevance map from the row associated with the CLS token in the output matrix, denoted by $\rho_{\text{CLS}}\in\mathbb{R}^{L}$, which contains the relevance scores evaluating each token’s influence on the classification token. Finally, to obtain the final explanation heatmap we reshape $\rho_{\text{CLS}}\in\mathbb{R}^{L}$ to $\sqrt{L}\times\sqrt{L}$ and upsample it back to the size of the original image using bilinear interpolation.
Although Mamba models are causal by definition, resulting in causal hidden attention matrices, our method can be extended to a bidirectional setting in a straightforward manner. This adaptation involves modifying Eq. 20 so that $\tilde{\mathbb{\alpha}}^{\lambda,d}$ becomes the outcome of summing the (two) per-direction matrices of the $\lambda$-layer and the $d$-channel.